OpenStudy (anonymous):

The cost to ship each order from a mail-order catalog depends on the weight of the order. A 20-pound order costs 10$ to ship. A 40-pound order costs 15$ to ship. Write a equation in slope intercept form that can be used to find the cost, C, to ship an order weighing p pounds.

OpenStudy (tkhunny):

So, C(x) = mx + b, where "b" is a flat cost per package and "m" is a cost related to the weight. In terms of these definitions, we are given: C(20) = m(20) + b = $10 C(40) = m(40) + b = $15 Can you solve the simultaneous equations for m and b? Really, just making the definitions was the more important part. Learn to WRITE DOWN clear definition.

OpenStudy (anonymous):

ty